I am 32, my sister is 18. How old will I be when I am 3x's her age?
Answer by stanbon(48545)

I am 32, my sister is 18. How old will I be when I am 3x's her age?
------------------------------
Let the year difference be "x" yrs.
Equation:
my age then = 3(sister's age then)
(32+x) = 3(18+x)
32+x = 54 + 3x
2x = -22
x = -11
--------------
I will be 32-11 = 21
Sister will be 18-11 = 7
==========================
Cheers,
Stan H.